# MMVA: Multimodal Matching Based on Valence and Arousal across Images, Music, and Musical Captions

### v1

@@ Introduction
Integrating multiple modalities in machine learning, particularly image, audio, and text, presents a compelling yet complex challenge. This complexity stems from the inherent diversity and intricate relationships of these three modalities. AudioCLIP (Guzhov et al. 2022) has established foundational work in this domain, capturing cross-modal relationships between non-musical audio, texts, and images. However, prior research did not specifically address music’s interplay with associated visual and textual components.
@@
To further this exploration, we expand the existing Image-Music-Emotion-Matching-Net (IMEMNet) (Zhao et al. 2020) by incorporating musical captions aligned with their audio counterparts. This expansion not only enriches the dataset but also challenges the conventional approach in multimodal research, which typically relies on one-to-one matching pairs between different modalities. For example, in AudioCLIP, an image of a cat is precisely matched with the audio of a cat meowing and the text “cat.” In contrast, the IMEMNet dataset deviates from this conventional paradigm. Instead of direct correlations between modalities, it features image-music matching based on continuous matching scores derived from valence (emotional positivity) and arousal (emotional intensity).
Leveraging IMEMNet, this research navigates a more nuanced landscape where the degree of multimodal correlation is variable and quantifiable only through a matching score. While previous work (Zhao et al. 2020) using the IMEMNet dataset employed predefined multimodal pairs, the current study implements random sampling of multimodal pairs during training. This approach yields superior performance on the valence-arousal (VA) prediction task compared to existing works, while also demonstrating the applicability of valence and arousal to various zero-shot tasks previously unexplored. Extensive experiments validate the effectiveness of the proposed method. This research thus not only introduces a novel set of modalities into the multimodal learning paradigm but also broadens the scope for potential applications of these combined modalities.
@@
In multimodal learning, a tower encoder refers to a separate neural network architecture that processes data of a particular modality, as exemplified by the CLIP image and text encoder of CLIP (Radford et al. 2021). Tower-based multimodal learning offers several advantages. First, each tower can be optimized for its specific data type, leveraging architectures best suited to the unique characteristics of the particular data modality. Second, it allows for the independent scaling and improvement of each tower without necessitating a redesign of the entire model. Updates or changes can be made to one modality without impacting the others.
Previous research has demonstrated that the audio modality can be successfully integrated with other modalities in multimodal learning using tower encoders. Conventional audio-text multimodal research (Wu et al. 2023; Elizalde et al. 2023; Mei et al. 2022) requires precise one-to-one audio-text matching in the speech data. Likewise, training a three-tower framework (Guzhov et al. 2022) capable of processing audio, image, and text relies on precise multimodal matching, as shown in Figure 1.
@@
Image-music multimodal learning remains largely unexplored, in contrast to music-text multimodal learning. For music-text tasks, MuLan (Huang et al. 2022) trains a text encoder and a music encoder in a two-tower framework for music tagging and music-text retrieval. For music-image tasks, CDCML (Zhao et al. 2020) constructed IMEMNet, a dataset of image-music pairs with VA matching scores, and proposed a multimodal training framework for emotion-based image-music matching. Employing a one-to-one image-music matching dataset, (Nakatsuka, Hamasaki, and Goto 2023) jointly trains image and music tower encoders to retrieve music from album covers and vice versa.
@@
The biggest challenge to multimodal learning involving image, music, and musical caption is data scarcity. To address this, IMEMNET-C extends the original IMEMNet dataset (Zhao et al. 2020) by incorporating musical captions corresponding to the music data in IMEMNet.
@@
IMEMNet incorporates three distinct real-world image datasets: IAPS (Lang 2005), NAPS (Marchewka et al. 2013), and EMOTIC (Kosti et al. 2017). The International Affective Picture System (IAPS) comprises 1,182 images, each annotated using a 9-point Valence-Arousal-Dominance (VAD) scale. The Nencki Affective Picture System (NAPS) includes 1,356 images evaluated on a 9-point bipolar semantic sliding scale, focusing on VA dimensions and approach-avoidance measures. The EMOTions In Context (EMOTIC) features 23,082 images of people, with annotations on a continuous 10-point VAD scale.
@@
IMEMNet utilizes the DEAM (Database for Emotion Analysis using Music) dataset, which consists of 1,802 music tracks, each annotated with valence and arousal values ranging from -1 to +1. These annotations are provided on a per-second basis and for the entire track. To ensure annotation stability, the annotation for each track starts at the 15th second. The tracks, predominantly 45 seconds in length and recorded at a frequency of 44100Hz, vary in duration, with some extending over 600 seconds. By dividing these tracks into two-second segments, the dataset comprises a total of 23,944 clips. Unlike Zhao et al. (2020), we extend the maximum clip length to 10 seconds by concatenating up to five consecutive clips with each two-second clip, allowing our framework to handle longer music data.
@@
It’s important to highlight that the potential number of all matching pairs in this setup is $a\times b$. Given the scale of the image and music clip datasets, this could result in hundreds of millions of possible pairs. To manage this vast number and prevent an overwhelming expansion of the dataset, Zhao et al. (2020) chose 50 images for each music clip: 30 are randomly picked from the image dataset, and the remaining 20 are split equally between those with the highest and lowest matching scores. To further refine the dataset, 10% of these pairs are randomly sampled to form the final IMEMNet dataset. Zhao et al. (2020) also ensured no overlap between the training, validation, and test sets.
@@
The IMEMNet dataset only contains image and music pairs, lacking musical captions for the music data. To address this gap, we introduce a new dataset, IMEMNet-C, which contains musical captions corresponding to the music data. Music captions are generated by the music-to-text large language model, LP-MusicCaps (Doh et al. 2023a).
Evaluations of the generated captions revealed the prevalence of redundant phrases, such as “The low quality recording.” To ensure high-quality captions for the music-image pairs, redundant phrases were removed in two steps. First, a rule-based refinement was applied to part of the audio captions. About half of the captions with redundant phrases contained the exact phrase “low quality recording features.” This redundant phrase was rephrased as “recording features.” Then, Llama-3.1-8B (Dubey et al. 2024) refined musical captions by removing and rephrasing the redundant expressions. We use the prompt ”Refine or remove an audio quality-related phrase to eliminate any mention of audio quality” with 4-shot examples.
@@
The AdamW (Loshchilov and Hutter 2017) optimizer is utilized with a learning rate of $0.0003$, a weight decay of $0.01$, $\beta_{1}$ of $0.9$, and $\beta_{2}$ of $0.999$. The training is conducted for 1000 epochs, using a batch size of 128 and a cosine annealing schedule (Loshchilov and Hutter 2016).
@@
For the image encoder $f_{img}$, MMVA employs the CLIP image encoder (Radford et al. 2021), which is based on ViT-B/32 (Dosovitskiy et al. 2021). $\boldsymbol{z}^{Img}$ corresponds to the $512$-dimensional class token vector. For both music and captions, MMVA uses BERT-based models (Devlin et al. 2019): MERT-95M (Li et al. 2023) for $f_{mus}$ and RoBERTa-base (Liu et al. 2019) for $f_{cap}$. For each of these BERT-based models, the proposed method extracts class tokens vectors $\boldsymbol{h}_{l}\in\mathbb{R}^{768}$ from the last $13$ transformer blocks (Vaswani et al. 2017) and gathers the vectors to form modality features $\boldsymbol{h}\in\mathbb{R}^{13\times 768}$. Subsequently, a 1D convolution is applied to derive a $768$-dimensional modality vector, which is then linearly projected to obtain the modality-specific embedding $\boldsymbol{z}^{k}\in\mathbb{R}^{512}$.
@@
We compare MMVA against the baseline results obtained by Zhao et al. (2020): SP-Net, $L^{3}$-Net (Arandjelović and Zisserman 2017), ACP-Net (Verma, Dhekane, and Guha 2019), and CDCML (Zhao et al. 2020). All the baselines are tailored for emotion-based image-music correspondence, employing the same emotion predictor architectures as used in our work. Evaluation employs Mean Squared Error (MSE) and Mean Absolute Error (MAE) as metrics to quantify the similarity between the predicted and ground-truth emotion values.
@@ Results.
Class-bin (Choi, Fazekas, and Sandler 2016) 4.0 13.8 20.1 8.3
Triplet-Glove (Choi et al. 2019) 2.8 11.2 18.6 6.6
Triplet-BERT (Won et al. 2021) 6.7 23.6 36.6 14.1
Contrastive-BERT (Doh et al. 2023b) 10.2 29.8 42.8 18.7
@@
Leveraging the VA-based similarity score (Equation 1), MMVA can achieve cross-modal retrieval for various modality pairs. Figure 3 illustrates the text-to-music retrieval based on valence and arousal. Given a query text, the music clip with the highest similarity score is retrieved. Table 2 demonstrates that MMVA achieves competitive retrieval performance. It is important to note that the baseline methods retrieve music clips from the Million Song Dataset (Bertin-Mahieux et al. 2011), which provides only pre-extracted audio features. In contrast, the MMVA music encoder operates directly on raw audio, making direct comparisons challenging due to differences in input representation and data.
@@
Figure 4 illustrates the image-to-text generation processes. First, a VA vector for the target image is obtained. Then, text prompts are exhaustively generated by inserting candidates into each text template illustrated in Figure 5. As a result, VA values for the generated text prompts are obtained, allowing for the calculation of the VA distance between the target image and the text prompts. By selecting the text prompt with the smallest distance, the best text prompt for the image is identified and fed into the text-to-music generative model, e.g. MusicGen (Copet et al. 2024). Generated captions are illustrated in Figure 6.
@@
To evaluate the effectiveness of arousal-based video summarization, we use the TVSum(Song et al. 2015) dataset, with frame-level importance scores assigned every two seconds. These scores range from 1 (not important) to 5 (very important). MMVA’s zeroshot approach is compared against well-established unsupervised video summarization methods: DR-DSN(Zhou and Qiao 2017), CSNet (Jung et al. 2018), AC-SUM-GAN (Apostolidis et al. 2021), and CA-SUM (Apostolidis et al. 2022). For generating video summaries, MMVA follows the same procedure as these baselines, solving a knapsack problem to select frames based on the frame-level importance scores, with each frame representing a short clip in a video. The duration of the summarized videos remained within 15% of the original.
@@
$\text{FS}=\frac{2(P\times R)}{P+R},P=\frac{||\hat{Q}\cap Q||}{||\hat{Q}||},R=\frac{||\hat{Q}\cap Q||}{||Q||}$ (5)
such that $P$ represents precision, and $R$ represents recall. Here, |||||$\cdot$||| | ⋅ | | denotes duration and $\cap$ signifies temporal overlap. Unlike the baseline models trained on 90% of the dataset and tested on the remaining 10%, MMVA is evaluated on the entire dataset. Generating video summaries with MMVA is zeroshot, eliminating the need for a train/test split.
@@ Acknowledgments
This work is supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) [RS-2024-00421203, RS-2024-00406127, 2021R1A2C3010887], the Ministry of Culture, Sports and Tourism & Korea Creative Content Agency [RS-2024-00399433], and Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City.

### v2

@@ Introduction
Integrating multiple modalities in machine learning, particularly image, audio, and text, presents a compelling yet complex challenge. This complexity stems from the inherent diversity and intricate relationships of these three modalities. AudioCLIP (Audioclip9747631) has established foundational work in this domain, capturing cross-modal relationships between non-musical audio, texts, and images. However, prior research did not specifically address music’s interplay with associated visual and textual components.
@@
To further this exploration, we expand the existing Image-Music-Emotion-Matching-Net (IMEMNet) (CDCML) by incorporating musical captions aligned with their audio counterparts. This expansion not only enriches the dataset but also challenges the conventional approach in multimodal research, which typically relies on one-to-one matching pairs between different modalities. For example, in AudioCLIP, an image of a cat is precisely matched with the audio of a cat meowing and the text “cat.” In contrast, the IMEMNet dataset deviates from this conventional paradigm. Instead of direct correlations between modalities, it features image-music matching based on continuous matching scores derived from valence (emotional positivity) and arousal (emotional intensity).
Leveraging IMEMNet, this research navigates a more nuanced landscape where the degree of multimodal correlation is variable and quantifiable only through a matching score. While previous work (CDCML) using the IMEMNet dataset employed predefined multimodal pairs, the current study implements random sampling of multimodal pairs during training. This approach yields superior performance on the valence-arousal (VA) prediction task compared to existing works, while also demonstrating the applicability of valence and arousal to various zero-shot tasks previously unexplored. Extensive experiments validate the effectiveness of the proposed method. This research thus not only introduces a novel set of modalities into the multimodal learning paradigm but also broadens the scope for potential applications of these combined modalities.
@@
In multimodal learning, a tower encoder refers to a separate neural network architecture that processes data of a particular modality, as exemplified by the CLIP image and text encoder of CLIP (CLIP). Tower-based multimodal learning offers several advantages. First, each tower can be optimized for its specific data type, leveraging architectures best suited to the unique characteristics of the particular data modality. Second, it allows for the independent scaling and improvement of each tower without necessitating a redesign of the entire model. Updates or changes can be made to one modality without impacting the others.
Previous research has demonstrated that the audio modality can be successfully integrated with other modalities in multimodal learning using tower encoders. Conventional audio-text multimodal research (wu2023large; CLAP2022; Mei2022metric) requires precise one-to-one audio-text matching in the speech data. Likewise, training a three-tower framework (Audioclip9747631) capable of processing audio, image, and text relies on precise multimodal matching, as shown in Figure 1.
@@
Image-music multimodal learning remains largely unexplored, in contrast to music-text multimodal learning. For music-text tasks, MuLan (huang2022mulan) trains a text encoder and a music encoder in a two-tower framework for music tagging and music-text retrieval. For music-image tasks, CDCML (CDCML) constructed IMEMNet, a dataset of image-music pairs with VA matching scores, and proposed a multimodal training framework for emotion-based image-music matching. Employing a one-to-one image-music matching dataset, (nakatsuka2023content) jointly trains image and music tower encoders to retrieve music from album covers and vice versa.
@@
The biggest challenge to multimodal learning involving image, music, and musical caption is data scarcity. To address this, IMEMNET-C extends the original IMEMNet dataset (CDCML) by incorporating musical captions corresponding to the music data in IMEMNet.
@@
IMEMNet incorporates three distinct real-world image datasets: IAPS (IAPS), NAPS (NAPS), and EMOTIC (EMOTIC). The International Affective Picture System (IAPS) comprises 1,182 images, each annotated using a 9-point Valence-Arousal-Dominance (VAD) scale. The Nencki Affective Picture System (NAPS) includes 1,356 images evaluated on a 9-point bipolar semantic sliding scale, focusing on VA dimensions and approach-avoidance measures. The EMOTions In Context (EMOTIC) features 23,082 images of people, with annotations on a continuous 10-point VAD scale.
@@
IMEMNet utilizes the DEAM (Database for Emotion Analysis using Music) dataset, which consists of 1,802 music tracks, each annotated with valence and arousal values ranging from -1 to +1. These annotations are provided on a per-second basis and for the entire track. To ensure annotation stability, the annotation for each track starts at the 15th second. The tracks, predominantly 45 seconds in length and recorded at a frequency of 44100Hz, vary in duration, with some extending over 600 seconds. By dividing these tracks into two-second segments, the dataset comprises a total of 23,944 clips. Unlike CDCML, we extend the maximum clip length to 10 seconds by concatenating up to five consecutive clips with each two-second clip, allowing our framework to handle longer music data.
@@
It’s important to highlight that the potential number of all matching pairs in this setup is $a\times b$. Given the scale of the image and music clip datasets, this could result in hundreds of millions of possible pairs. To manage this vast number and prevent an overwhelming expansion of the dataset, CDCML chose 50 images for each music clip: 30 are randomly picked from the image dataset, and the remaining 20 are split equally between those with the highest and lowest matching scores. To further refine the dataset, 10% of these pairs are randomly sampled to form the final IMEMNet dataset. CDCML also ensured no overlap between the training, validation, and test sets.
@@
The IMEMNet dataset only contains image and music pairs, lacking musical captions for the music data. To address this gap, we introduce a new dataset, IMEMNet-C, which contains musical captions corresponding to the music data. Music captions are generated by the music-to-text large language model, LP-MusicCaps (musiccaps).
Evaluations of the generated captions revealed the prevalence of redundant phrases, such as “The low quality recording.” To ensure high-quality captions for the music-image pairs, redundant phrases were removed in two steps. First, a rule-based refinement was applied to part of the audio captions. About half of the captions with redundant phrases contained the exact phrase “low quality recording features.” This redundant phrase was rephrased as “recording features.” Then, Llama-3.1-8B (LLAMA3) refined musical captions by removing and rephrasing the redundant expressions. We use the prompt ”Refine or remove an audio quality-related phrase to eliminate any mention of audio quality” with 4-shot examples.
@@
The AdamW (adamw) optimizer is utilized with a learning rate of $0.0003$, a weight decay of $0.01$, $\beta_{1}$ of $0.9$, and $\beta_{2}$ of $0.999$. The training is conducted for 1000 epochs, using a batch size of 128 and a cosine annealing schedule (cosanneal).
@@
For the image encoder $f_{img}$, MMVA employs the CLIP image encoder (CLIP), which is based on ViT-B/32 (vit). $\boldsymbol{z}^{Img}$ corresponds to the $512$-dimensional class token vector. For both music and captions, MMVA uses BERT-based models (BERT): MERT-95M (MERT) for $f_{mus}$ and RoBERTa-base (roberta) for $f_{cap}$. For each of these BERT-based models, the proposed method extracts class tokens vectors $\boldsymbol{h}_{l}\in\mathbb{R}^{768}$ from the last $13$ transformer blocks (transformer) and gathers the vectors to form modality features $\boldsymbol{h}\in\mathbb{R}^{13\times 768}$. Subsequently, a 1D convolution is applied to derive a $768$-dimensional modality vector, which is then linearly projected to obtain the modality-specific embedding $\boldsymbol{z}^{k}\in\mathbb{R}^{512}$.
@@
We compare MMVA against the baseline results obtained by CDCML: SP-Net, $L^{3}$-Net (L3-NET), ACP-Net (ACP-Net), and CDCML (CDCML). All the baselines are tailored for emotion-based image-music correspondence, employing the same emotion predictor architectures as used in our work. Evaluation employs Mean Squared Error (MSE) and Mean Absolute Error (MAE) as metrics to quantify the similarity between the predicted and ground-truth emotion values.
@@ Results.
Class-bin (classbin) 4.0 13.8 20.1 8.3
Triplet-Glove (triplet-glove) 2.8 11.2 18.6 6.6
Triplet-BERT (triplet-bert) 6.7 23.6 36.6 14.1
Contrastive-BERT (contrastive-bert) 10.2 29.8 42.8 18.7
@@
Leveraging the VA-based similarity score (Equation 1), MMVA can achieve cross-modal retrieval for various modality pairs. Figure 3 illustrates the text-to-music retrieval based on valence and arousal. Given a query text, the music clip with the highest similarity score is retrieved. Table 2 demonstrates that MMVA achieves competitive retrieval performance. It is important to note that the baseline methods retrieve music clips from the Million Song Dataset (millionsong), which provides only pre-extracted audio features. In contrast, the MMVA music encoder operates directly on raw audio, making direct comparisons challenging due to differences in input representation and data.
@@
Figure 4 illustrates the image-to-text generation processes. First, a VA vector for the target image is obtained. Then, text prompts are exhaustively generated by inserting candidates into each text template illustrated in Figure 5. As a result, VA values for the generated text prompts are obtained, allowing for the calculation of the VA distance between the target image and the text prompts. By selecting the text prompt with the smallest distance, the best text prompt for the image is identified and fed into the text-to-music generative model, e.g. MusicGen (musicgen). Generated captions are illustrated in Figure 6.
@@
To evaluate the effectiveness of arousal-based video summarization, we use the TVSum(TVSUM) dataset, with frame-level importance scores assigned every two seconds. These scores range from 1 (not important) to 5 (very important). MMVA’s zeroshot approach is compared against well-established unsupervised video summarization methods: DR-DSN(DR-DSN), CSNet (CSNet), AC-SUM-GAN (AC-SUM-GAN), and CA-SUM (CA-SUM). For generating video summaries, MMVA follows the same procedure as these baselines, solving a knapsack problem to select frames based on the frame-level importance scores, with each frame representing a short clip in a video. The duration of the summarized videos remained within 15% of the original.
@@
[…]
such that $P$ represents precision, and $R$ represents recall. Here, $||\cdot||$ denotes duration and $\cap$ signifies temporal overlap. Unlike the baseline models trained on 90% of the dataset and tested on the remaining 10%, MMVA is evaluated on the entire dataset. Generating video summaries with MMVA is zeroshot, eliminating the need for a train/test split.
@@ Acknowledgments
This work is supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) [RS-2024-00421203, RS-2024-00406127, 2021R1A2C3010887], Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2024 (Project Name: Development of K-POP artist-centered video editing solution: customized multimodal AI model and generative asset, Project Number: RS-2024-00399433, Contribution Rate: 50%), and Artificial intelligence industrial convergence cluster development project funded by the Ministry of Science and ICT (MSIT, Korea) & Gwangju Metropolitan City.